What value of x will make the equation 3(x + 15) - 6x = -6(x - 3) true?
Answer by stanbon(75887)

What value of x will make the equation 3(x + 15) - 6x = -6(x - 3) true?
-------
3x + 45 - 6x = -6x + 18
------------------------------
3x = -27
x = -9
==============
Cheers,
Stan H.
--------------